Achondroplastic dwarfism is an inherited condition. In one study, 108 achondroplastic dwarfs produced 27 offspring, while 457 non-dwarf individuals produced 582 children. What is the relative fitness of the achondroplastic dwarf genotype?

Please explain your answer.
Answer #1

Relative fitness is the ratio of the offspring produced by any each individual of a genotype to the number of offspring produced by the genotype with the highest number of offspring per individual.

Here,

Number of individuals with Achondroplastic dwarfism = 108

Number of Non-dwarf individuals = 457

Number of offspring of non-dwarf individuals = 582

Number of offspring of Achondroplastic dwarf individuals = 27

Reproductive rate of Achondroplastic dwarf individuals
   = Number of offspring of Achondroplastic dwarf individuals/Number of individuals with Achondroplastic dwarfism
= 27/108 = 0.25

Reproductive rate of non-dwarf individuals
= Number of offspring of non-dwarf individuals/Number of non-dwarf individuals
= 582/457 = 1.27

Therefore relative fitness of the Achondroplastic dwarf genotype
= Reproductive rate of Achondroplastic dwarf individuals/Reproductive rate of non-dwarf individuals
= 0.25/1.27 = 0.196